OpenStudy (anonymous):

A triangular prism has a volume of 5 m3. If each dimension of the prism is tripled, what is the new volume of the prism? Answer 135 m3 45 m3 15 m3 405 m3

OpenStudy (anonymous):

and i chose 15

OpenStudy (anonymous):

well since Volume of a triangular prism is V=5m^3 and the formula is: \[\LARGE V=\frac12 bhl\] now , if EACH dimension of the prism is tripled which mean 3times... we have \[\LARGE V=\frac12 \cdot 3\cdot b\cdot 3\cdot h\cdot 3\cdot l\] so in total there is 81 times Volume of the Prism we had at principal... which means if \[\LARGE V=\frac12 bhl\] \[\LARGE V=5 m^3\] and now \[\LARGE V=81\cdot \frac12 bhl\] we also raise 5m^3 by \[\LARGE V=81\cdot 5 m^3\] and when we multiply it out, we see that the new Volume of this prism is: V=405 m^3